If X and Y are independent Poisson variates such that P(X = 1)=P( X =2) and P(Y=2)= P(Y= 3).
Find the variance of X − 2Y .
1
Expert's answer
2020-03-09T14:16:42-0400

If X and Y are independent Poisson variables such that P(X = 1)=P( X =2) and P(Y=2)= P(Y= 3).

Find the variance of X − 2Y .


"P(X=x)={e^{-\lambda_1}\lambda_1^x \over x!}"

Given that "P(X=1)=P(X=2)." Then


"{e^{-\lambda_1}\lambda_1^1 \over 1!}={e^{-\lambda_1}\lambda_1^2 \over 2!}, \lambda_1>0""\lambda_1=2"

"P(Y=y)={e^{-\lambda_2}\lambda_2^x \over x!}"

Given that "P(Y=2)=P(Y=3)." Then


"{e^{-\lambda_2}\lambda_2^2 \over 2!}={e^{-\lambda_2}\lambda_2^3 \over 3!}, \lambda_2>0""{\lambda_2^2 \over 2}={\lambda_2^3 \over6}""\lambda_2=3."

For a Poisson random variable "Z, Var(Z)=\lambda." Then


"Var(X)=\lambda_1=2, \ Var(Y)=\lambda_2=3"

Hence


"Var(X-2Y)=Var(X)+2^2Var(Y)=""=2+4(3)=14"

"Var(X-2Y)=14"